find an equation of the line that has slope m and y-intercept b. m=-2, b= -1
Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
m = -2 and b = -1
Line
y = (-2)x + (-1)
y = -2x - 1 or 2x + y = -1 or 2x + y + 1 = 0